Please help with title: solving systems by elimination
A chemist has a beaker of a 3% acid solution and a beaker of a 7% acid solution. He needs to make 75ml of a 4% acid solution. complete the table.
3% solutiom + 7% solution = 4% solution
Amt of solution x + y =
Amt of acid (ml) ____x + ____y =

Answer by checkley77(12844)

.07X+.03(75-X)=.04*75
.07X+2.25-.03X=3
.04X=3-2.25
.04X=.75
X=.75/.04
X=18.75 ml. OF 7% SOLUTION.
75-18.75=56.25 ml. OF 3% SOLUTION.
PROOF:
.07*18.75+.03*56.25=.04*75
1.3125+16.875=3
3=3